### v1

@@ 4 Discussion
In this work, we have introduced MERGE-RNA, a physics-based framework for the analysis of chemical probing data that addresses fundamental limitations of conventional approaches. Our approach moves beyond the use of experimental data as simple pseudo-energetic restraints [deiganAccurateSHAPEdirectedRNA2009] and instead models the entire experimental pipeline, from the physical interaction of the probe with the RNA to the final mutational readout. This integrated approach allows for the determination of a set of transferable, physically meaningful parameters that describe the experimental pipeline, and ultimately enables the integration of multiple datasets for the inference of the unperturbed, native structural ensemble. As demonstrated by our successful deconvolution of mixed structural states (both on synthetic and experimental data), our method provides a powerful tool for quantitatively characterizing these populations. This capability is crucial for understanding the mechanisms of functional RNAs and for dissecting complex processes such as strand displacement and co-transcriptional folding, where the molecule transiently adopts a series of alternative conformations.
A central strength of our methodology is its ability to avoid over-constraining structural predictions. Standard approaches that convert reactivity data into pseudo-free energies tend to modify the energy landscape globally, often altering regions that are already correctly predicted. As previously noted, an ideal method should refrain from changing predictions that are already accurate and instead focus on introducing only the minimal adjustments needed to reconcile the model with the data, thereby highlighting where the baseline model is insufficient. Our framework achieves this by applying the maximum entropy principle to infer the minimal soft constraints ($\lambda_{i}$) required to fit the data. This ensures that modifications arise only where the data provide strong evidence of a deviation from the initial model, refining the energy landscape locally while preserving the predictive power of the underlying thermodynamics. This manner of correcting the ensemble resembles the one introduced in Washielt et al [washietlRNAFoldingSoft2012] to analyze SHAPE data, although the derivation is different, and it improves in the capability of modeling the probe-RNA interactions directly.
The optimization of the structural ensemble is here done using the maximum entropy principle, which, among the ensembles with highest compatibility with experiment, choose the one that is as close as possible to the prior. The variability of the result was here limited by adding boundaries on the resulting soft constraints. Alternative approaches based on the use of entropy as a regularization term modulated by a chosen hyperparameter might be explored [hummerBayesianEnsembleRefinement2015, cesariUsingMaximumEntropy2018].
@@
Existing ensemble-deconvolution pipelines such as DREEM [tomezskoDeterminationRNAStructural2020], DRACO [morandiGenomescaleDeconvolutionRNA2021], and DANCE-MaP [olsonDiscoveryLargescaleCellstateresponsive2022] cluster sequencing reads with multiple co-occurring modifications and assign each cluster to a discrete structure. By capitalizing on per-read correlation patterns, these methods recover information about alternative conformations but at the cost of discarding the abundant reads that contain zero or single modifications and demanding deep coverage to resolve low-probability states. Our framework instead calibrates a physical model against the averaged mutational fractions, ensuring that every read (even negative ones) contributes signal even in sparse-modification regimes. The trade-off is the loss of explicit per-read co-variation, yet in principle the explicit reads can be integrated into our framework to fine-tune the populations of specific structures. We leave this as a possible future extension of the model. Moreover, read-level clustering can struggle to distinguish highly similar conformations and to represent continuous ensembles comprising many closely related intermediates (e.g., along strand-displacement pathways), which gives our approach an advantage. Consequently, our framework is well-suited not only to capture discrete alternative structures but also to quantify continuous ensembles that enable interpretation of dynamic RNA processes. Furthermore, alternative methods implicitly assume that each read originates from a single structure, which may not hold for RNAs undergoing rapid conformational dynamics and/or long probing timescales. Our framework instead models the ensemble-averaged behavior directly, naturally accommodating fast interconversions between states at equilibrium. Strikingly, even without using individual reads, the physics of the model enables (a) reconstructing correlations (e.g., the cooperative formation of two distant loops) and (b) measuring the population of minor substates (e.g., the strand-displaced intermediates).
A key feature of our calibrated framework is the ability to predict the structural ensemble at any probe concentration, and most importantly, to extrapolate to zero probe concentration to obtain the native state. In the specific cases analyzed in this study, the dependence of the base-pairing probabilities on probe concentration was relatively modest, which introduces partial under-determination in the estimated parameters (e.g., increasing $\mu$ can mimic decreasing $p_{\mathrm{bind}}$). This does not compromise robustness or transferability of the calibrated parameters (as shown above e.g. in Fig.2a) and can be mitigated by widening the concentration range. As a future direction, a concise Bayesian treatment (priors + posterior curvature) could quantify identifiability and provide credible intervals. As perspective, one may place weakly informative priors and consider $\mathbb{P}(\theta\mid\text{data})$; Maximum a posteriori would remain a practical point estimate, while posterior would yield credible intervals and identifiability diagnostics. However, the model is fully equipped to handle systems where the probe induces significant non-linearities [calonaciMolecularDynamicsSimulations2023] rising form structural rearrangement, denaturation, or saturation. This behavior normally presents challenges in chemical probing experiments and suggests that one should work in the single-hit-kinetics regime [arnoldCooperativityRNAChemical2023], which is however difficult for large RNAs and effectively eliminates correlation information. The ability to explicitly model and remove these concentration-dependent perturbations is a significant step towards obtaining a more faithful representation of an RNA’s native conformational landscape.
The explicit separation of intrinsic RNA thermodynamics (captured by the baseline model and the $\lambda_{i}$ soft constraints) from the probe-specific interaction physics (captured by parameters such as $\mu_{r}$ and $\Delta\mu_{\text{pairing}}$) opens a natural avenue for future extensions. One could readily extend this framework to simultaneously analyze data from multiple, distinct chemical probes such as CMCT [ehresmannProbingStructureRNAs1987], SHAPE [merinoRNAStructureAnalysis2005, weeksAdvancesRNAStructure2010, marinusNovelSHAPEReagent2021], hydroxyl radical probing [adilakshmiHydroxylRadicalFootprinting2006], and ETS [doudsNewReagentVivo2024]. IPANEMAP [saaidiIPANEMAPIntegrativeProbing2020] has shown that an improve on performance is obtainable by integrating multi-probe measurements, but relies on empirical reactivity-to-energy conversions. In our framework, each probe would be described by its own set of physical parameters, but all datasets would be used to constrain and refine a single, shared set of $\lambda_{i}$ parameters. As an example, SHAPE reagents could be modeled with the same formalism introduced here for DMS, but with different binding preferences (eight new $p_{\text{bind}}$, two for each nucleotide) and energetic penalties (new $\mu_{r}$ and $\Delta\mu_{\text{pairing}}$). This would yield a more robust and comprehensive determination of the structural ensemble, leveraging the orthogonal chemical information provided by different reagents and enabling the integration of multiple and more diverse datasets.
@@
An additional extension concerns interactions with other biomolecules (e.g., RNA-protein binding) that can confound probing readouts. If reliable protein-RNA contacts are known on specific segments, those positions could be excluded from the loss to avoid bias from probe occlusion and simultaneously constrained to avoid internal pairing. Conversely, when interaction sites are unknown, the calibrated model could be used to localize discrepancies between predicted and observed mutational profiles, flagging sequence regions likely involved in protein or ligand interactions for follow-up experiments.
@@
We acknowledge several important limitations of the current model. The underlying ViennaRNA framework does not account for the formation of pseudoknots, which are known to be critical for the function of many RNAs. Even in absence of pseudoknots, knowledge of the secondary structure alone might not be sufficient to predict reactivities towards DMS probing, which might dependent on 3D structural features [deenalatthaCharacterizing3DRNA2025]. Furthermore, while we have shown that the physical parameters are transferable across different RNA molecules under constant experimental conditions, their transferability between experiments conducted under varying conditions (e.g., different temperatures, buffers, or duration of exposition to the probe) is not guaranteed, therefore, we recommend fitting data from experiments conducted under similar conditions as a cohesive group. To achieve broader applicability, the model would need to be extended to explicitly describe the dependence of the physical parameters on these experimental variables.
@@ 5 Data and code availability
The MERGE-RNA source code, together with analysis scripts, is available at https://github.com/giusSacco/MERGE-RNA.
@@ 8 Acknowledgments
We thank A. Mustoe and J. D. Yesselman for useful discussions.
{appendices}
9 Supplementary information
9.1 Physical Model
The physical model underlying our analysis of RNA chemical probing experiments is illustrated in the left panel of Figure 1 and will be detailed in this section. We derive and justify the expression for the free energy of a secondary structure $s$ in the presence of a chemical probe (as reported in the main text Eq. 3) which enables the prediction of the ensemble of structures, and explain our methodology for producing a prediction of the reactivity profiles at varying probe concentrations, which serve as the theoretical basis for fitting experimental data.
@@ 9.1 Physical Model
where $N\equiv N_{u}(s)+N_{p}(s)$ is the sequence length. The first term is structure-independent and does not contribute to the populations in the experiment, so it can be safely ignored in our calculations. In practice, this formulation enables us to account for the probe influence on the secondary structure populations by adding $\ln\left(\frac{1+\mathrm{e}^{\beta\mu^{\prime}}}{1+\mathrm{e}^{\beta\mu}}\right)$ as a soft constraint in ViennaRNA for each paired bases, allowing us to efficiently compute pairing probabilities within its computational framework. By substituting the expression for $\Delta F_{\text{[DMS]}}(s)$ into Equation 7, we obtain the final expression for the free energy of a secondary structure
@@
9.1.1 Expected Mutation Rate and Systematic Biases
@@
Worth mentioning in this summary–altough not fitted–are the $\epsilon_{i}$, which are calibrated from the control experiment and incorporate site-specific systematic errors shared across replicates and concentrations, and effectively serve as offsets to the predicted mutation profiles. They enable the decoupling of experimental biases from the thermodynamic and concentration-dependent effects, allowing the model to focus on physically relevant signal.
9.2 Model Training
9.2.1 Loss function
@@ 9.2.2 Fit of parameters
The optimization is performed in a two-stage process to minimize the loss function. In the first stage, we fit the physical parameters that are common to all experiments, keeping the sequence-specific soft constraints ($\lambda_{i}$) fixed. This global optimization starts from random initializations and runs until convergence. In the second stage, the optimized physical parameters are held constant while we fit the $\lambda_{i}$ soft constraints independently for each RNA sequence. This hierarchical strategy allows us to first establish a robust physical model and then refine the seondary structure populations with sequence-specific adjustments. During the optimization, the $\lambda_{i}$ parameters are constrained to the range [1,1][-1,1] kcal/mol to ensure proper regularization and avoid overfitting.
9.3 Experimental Procedures
@@
9.4 Mapping and counting of mutations
@@
Supplementary Information
Training sets $\mu_{r}$ $\Delta\mu_{\text{pairing}}$ $p_{\text{bind}}(\text{A},\text{unpaired})$ $p_{\text{bind}}(\text{C},\text{unpaired})$ $p_{\text{bind}}(\text{G})$ $p_{\text{bind}}(\text{U})$ $m_{1}$
@@
Table 1: Physical parameters obtained by training on different triplets of RNA systems.
Figure 6:
Figure 7:
Figure 8:
Figure 9:
Figure 10:
Figure 11:
Figure 12:
Figure 13:
Table 2: Loop co-occupancy inferred from baseline thermodynamic ensembles and after MERGE-RNA refinement. Entries report the percentage of structures in which Loop1 and Loop2 are formed.

### v2

@@ Discussion
In this work, we have introduced MERGE-RNA, a physics-based framework for the analysis of chemical probing data that addresses fundamental limitations of conventional approaches. Our approach moves beyond the use of experimental data as simple pseudo-energetic restraints [deiganAccurateSHAPEdirectedRNA2009] and instead models the entire experimental pipeline, from the physical interaction of the probe with the RNA to the final mutational readout. This enables the determination of transferable, physically meaningful parameters and the integration of multiple datasets for inferring the native structural ensemble. As demonstrated by our successful deconvolution of mixed structural states (both on synthetic and experimental data), our method provides a powerful tool for quantitatively characterizing conformational populations, crucial for understanding functional RNAs and dissecting processes such as strand displacement and co-transcriptional folding.
We note that our model explicitly includes all four nucleotides even though DMS exhibits structure-dependent reactivity primarily with A and C. Including the complete sequence leads to more robust estimates of the physical parameters, ultimately improving the accuracy of inferred structural ensembles.
A central strength is the ability to avoid over-constraining: standard pseudo-free energy approaches [deiganAccurateSHAPEdirectedRNA2009] modify the energy landscape globally, often altering already-correct regions. Similarly to Ref. [washietlRNAFoldingSoft2012], our framework infers minimal soft constraints, ensuring modifications arise only where data support deviation, and resulting in a lower loss (Fig. 2b). In addition, our physics-based model naturally enables the combination of multiple datasets in a single fitting process. Among ensembles compatible with experiment, the maximum entropy principle chooses the one closest to the prior, and modifications of the ensemble are here regularized by bounding soft constraints. Alternative regularization approaches might be explored [hummerBayesianEnsembleRefinement2015, cesariUsingMaximumEntropy2018].
@@
Existing ensemble-deconvolution pipelines such as DREEM [tomezskoDeterminationRNAStructural2020], DRACO [morandiGenomescaleDeconvolutionRNA2021], and DANCE-MaP [olsonDiscoveryLargescaleCellstateresponsive2022] cluster sequencing reads with multiple co-occurring modifications and assign each cluster to a discrete structure. By capitalizing on per-read correlation patterns, these methods recover information about alternative conformations but at the cost of discarding the abundant reads with zero or single modifications and demanding deep coverage to resolve low-probability states. Alternative ensemble fitting approaches include Rsample [spasicModelingRNASecondary2018] and SLEQ [liStatisticalModelingRNA2018]. Rsample employs a sample-and-cluster approach, reweighting the Boltzmann ensemble through pseudo-free energy adjustments resembling our soft constraints, though without physics-grounded calibration, but rather adopting an empirical strategy in a non self-consistent fashion. SLEQ uses a sample-and-select strategy rooted in the maximum parsimony principle to identify the sparsest set of structures explaining modification patterns, differing from our maximum entropy approach designed to minimally perturb the thermodynamic model. Our framework calibrates against averaged mutational fractions, ensuring every read (even negatives) contributes signal in sparse-modification regimes. The trade-off is loss of explicit per-read co-variation, though reads could in principle be integrated to fine-tune populations—a possible future extension. Read-level clustering would struggle to distinguish multiple similar conformations and represent continuous ensembles with many intermediates (e.g., strand-displacement pathways). Furthermore, alternative methods assume each read originates from a single structure, which may not hold for RNAs undergoing rapid dynamics. Our framework models ensemble-averaged behavior directly, accommodating fast interconversions at equilibrium. Consequently, our framework captures both discrete alternative structures and continuous ensembles enabling interpretation of dynamic processes. Strikingly, even without individual reads, the model reconstructs correlations and measures minor substate populations, which would be missed in methods that enumerate a small number of representative structures.
A key feature is predicting the structural ensemble at any probe concentration, especially extrapolating to zero concentration for the native state. In our analyses, base-pairing probability dependence on concentration was modest, introducing partial parameter under-determination (e.g., increasing $\mu$ mimics decreasing $p_{\mathrm{bind}}$). This does not compromise robustness or transferability (Fig.2a) and can be mitigated by widening the concentration range; Bayesian treatment could be used to quantify identifiability. At the same time, the model is equipped to handle systems where probes induce significant non-linearities [calonaciMolecularDynamicsSimulations2023] from rearrangement, denaturation, or saturation. Standard protocols recommend avoiding these effects by reaching the single-hit-kinetics [arnoldCooperativityRNAChemical2023], which might be difficult or impossible for large RNAs, and intrinsically eliminates correlation information. Explicitly modeling these concentration-dependent perturbations might enable a more faithful representation of native conformational landscapes.
The separation of RNA thermodynamics (baseline model and $\lambda_{i}$) from probe-specific physics (all fo the other parameters) enables future multi-probe extensions. One could analyze data from CMCT [ehresmannProbingStructureRNAs1987], SHAPE [merinoRNAStructureAnalysis2005, weeksAdvancesRNAStructure2010, marinusNovelSHAPEReagent2021], hydroxyl radical footprinting [adilakshmiHydroxylRadicalFootprinting2006], and ETS [doudsNewReagentVivo2024] simultaneously. IPANEMAP [saaidiIPANEMAPIntegrativeProbing2020] showed improvements from multi-probe integration but relies on empirical conversions. In our framework, each probe would be modeled with its own parameters while each molecule would have specific $\lambda_{i}$ values, leveraging orthogonal chemical information.
@@
An additional extension concerns RNA-protein interactions that confound probing readouts. Known contacts could be excluded from the loss and constrained against internal pairing; unknown interaction sites could be flagged by localizing discrepancies between predicted and observed profiles.
We note that validation of RNA structure prediction methods typically relies on reference structures from crystallography, cryo-EM, or covariance analysis, each with inherent biases. Crystals often fail to form for conformationally heterogeneous molecules, selecting against dynamic ensembles. Cryo-EM reconstructions typically use only a small fraction of recorded particles, with difficult-to-estimate effects on captured dynamics. Covariance-based methods reflect evolutionarily conserved, biologically functional structures, which may represent only a subset of conformations observed in vitro. These considerations highlight the value of probing-based ensemble methods as a complementary approach to traditional structure determination.
@@
We acknowledge important limitations. The underlying ViennaRNA framework does not account for pseudoknots, critical for many RNAs. More generally, secondary structure alone may not suffice to predict DMS reactivities, which can depend on 3D features [deenalatthaCharacterizing3DRNA2025]. While physical parameters transfer across RNAs under constant conditions, transferability between varying conditions (temperature, buffer, probing duration) is not guaranteed; we recommend fitting similar-condition data together. Broader applicability would require modeling parameter dependence on experimental variables.
@@ 1 Data and code availability
The MERGE-RNA source code, together with analysis scripts, is available at https://github.com/giusSacco/MERGE-RNA. Sequencing data for all the experiments performed in this work have been deposited in the European Nucleotide Archive (ENA) under accession number XXXXXXXXX.
@@ 4 Acknowledgments
We thank Catherine Eichhorn, Anthony Mustoe and Joseph Yesselman for useful discussions. Anthony Mustoe is also acknowledged for reading a preliminary version of this manuscript and providing useful suggestions. RPS acknowledges the interdisciplinary Thematic Institute IMCBio+, as part of the ITI 2021–2028 program of the University of Strasbourg, CNRS and Inserm, IdEx Unistra (ANR-10-IDEX-0002), SFRI-STRAT’US (ANR 20-SFRI-0012), and EUR IMCBio (ANR-17-EURE-0023) under the framework of the French Investments of the France 2030 Program.
Supplementary Information
Physical Model
The physical model underlying our analysis of RNA chemical probing experiments is illustrated in the left panel of Figure 1 and will be detailed in this section. We derive and justify the expression for the free energy of a secondary structure $s$ in the presence of a chemical probe (as reported in the main text) which enables the prediction of the ensemble of structures, and explain our methodology for producing a prediction of the reactivity profiles at varying probe concentrations, which serve as the theoretical basis for fitting experimental data.
@@
where $N\equiv N_{u}(s)+N_{p}(s)$ is the sequence length. The first term is structure-independent and does not contribute to the populations in the experiment, so it can be safely ignored in our calculations. In practice, this formulation enables us to account for the probe influence on the secondary structure populations by adding $\ln\left(\frac{1+\mathrm{e}^{\beta\mu^{\prime}}}{1+\mathrm{e}^{\beta\mu}}\right)$ as a soft constraint in ViennaRNA for each paired base, allowing us to efficiently compute pairing probabilities within its computational framework. By substituting the expression for $\Delta F_{\text{[DMS]}}(s)$ into Equation 7, we obtain the final expression for the free energy of a secondary structure
@@
Expected Mutation Rate and Systematic Biases
@@
Worth mentioning in this summary–although not fitted–are the $\epsilon_{i}$, which are calibrated from the control experiment and incorporate site-specific systematic errors shared across replicates and concentrations, and effectively serve as offsets to the predicted mutation profiles. They enable the decoupling of experimental biases from the thermodynamic and concentration-dependent effects, allowing the model to focus on physically relevant signal.
Model Training
Loss function
@@
The optimization is performed in a two-stage process to minimize the loss function. In the first stage, we fit the physical parameters that are common to all experiments, keeping the sequence-specific soft constraints ($\lambda_{i}$) fixed. This global optimization starts from random initializations and runs until convergence. In the second stage, the optimized physical parameters are held constant while we fit the $\lambda_{i}$ soft constraints independently for each RNA sequence. This hierarchical strategy allows us to first establish a robust physical model and then refine the secondary structure populations with sequence-specific adjustments. During the optimization, the $\lambda_{i}$ parameters are constrained to the range [1,1][-1,1] kcal/mol to ensure proper regularization and avoid overfitting.
Experimental Procedures
@@
Mapping and counting of mutations
@@
Supplementary Figures and Tables
Training sets $\mu_{r}$(kcal/mol) $\Delta\mu_{\text{pairing}}$ (kcal/mol) $p_{\text{bind}}(\text{A},\text{unpaired})$ $p_{\text{bind}}(\text{C},\text{unpaired})$ $p_{\text{bind}}(\text{G})$ $p_{\text{bind}}(\text{U})$ $m_{1}$
@@
Table S1: Physical parameters obtained by training on different triplets of RNA systems.
Figure S1:
Figure S2:
Figure S3:
Figure S4:
Figure S5:
Figure S6:
Figure S7:
Figure S8:
Figure S9:
Figure S10:
Figure S11:
Figure S12:
Table S2: Loop co-occupancy inferred from baseline thermodynamic ensembles and after MERGE-RNA refinement. Entries report the percentage of structures in which Loop1 and Loop2 are formed.